# Inattention to States and Characteristics

### v1

@@ Introduction
Example: state-action with costlessly observed $\theta$. Consider a state-action model in which the DM knows their type $\theta$ a priori. The DM has an action set $A=\{1,...,J\}$ consisting of $J$ products with $J$ characteristics. The DM knows what the characteristics are, but does not know how they are permuted among the $J$ actions. Let $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$ denote these characteristics. Denote $\xi$ with the characteristic of the product he eventually chooses. The mutual information $I_{P}(j,\boldsymbol{\xi})$ between action $j\in[J]$ and $\boldsymbol{\xi}$ is
$\int D_{KL}(P(\boldsymbol{\xi}|j)\|P(\boldsymbol{\xi}))\>dP(j)=D_{KL}(P(\boldsymbol{\xi}|j=1)\|P(\boldsymbol{\xi}))=D_{KL}(P(\xi_{1}|j=1)\|P(\xi_{1}))=D_{KL}(P(\xi)\|\phi(\xi))$
@@
The third equality follows from the observation that choosing $j=1$ provides only information about what $\xi_{1}$ and is not (meaning it provides no information about $(\xi_{2},...,\xi_{J})$ beyond the fact that they are not $\xi_{1}$). Thus, this corresponds to the state-characteristic model with $\alpha=1$ and $\phi(\xi)=1/J$.
@@
Example: Maxwell-Boltzmann statistics. When $\alpha=1$, the law of iterated expectations yields
@@
$\frac{dP(\xi,\theta)}{d(P|_{\xi}\otimes\mu)(\xi,\theta)}=\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}<\infty$ $\frac{dP(\xi)}{d\phi(\xi)}<\infty$ $\frac{dP(\theta)}{d\mu(\theta)}=1$ (2)
@@
If $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$, then $P(\cdot|\theta)\ll\phi$ $\mu$-a.s. and $P\ll\phi\otimes\mu$.
@@
Let $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$. There exist valid densities such that eq.˜2 is satisfied.
@@
We consider the case where $\alpha\in(0,1)$. We normalize $\lambda$ to $1$, since maximizing $U(P)$ is the same as maximizing $U(P)/\lambda$. To obtain results for $\lambda\neq 1$, simply take all results herein in which $u(\xi,\theta)$ and associated additive terms appear and divide them by $\lambda$.
@@
If $P$ solves the objective eq.˜1 then
@@
If $P$ solves the objective eq.˜1 then
@@
The optimal marginal $P|_{\xi}$ is unique, and therefore the optimal CCP is unique.
@@
If $P$ maximizes $V$, then $P|_{\xi}\ll\phi\ll P|_{\xi}$.
These follow from corollaries˜7.1, 8 and 15.1. In the discrete case, these can be derived by taking Lagrangians.
@@
$\frac{P(\xi|\theta)}{P^{\prime}(\xi|\theta)}=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}}{\phi^{\prime}(\xi)^{\alpha}P^{\prime}(\xi)^{1-\alpha}}\frac{Z^{\prime}(\theta;P^{\prime}|_{\xi})}{Z(\theta;P|_{\xi})}$
@@
First-Step Orthogonality and Its Implications
First-step orthogonality is a property of moments of the form $\int f_{P}(\omega)\>dP(\omega)$, where the integrand $f$ depends on the probability with respect to which we are integrating. Loosely, it states that a perturbation in $P$ locally affects the moment only through the re-weighting $dP$. It constitutes a sufficient condition for the Gibbs property and additive separability. It is a relatively unique property of entropy-based costs which essentially comes from the fact that $\frac{d}{dx}\log x=1/x$.
@@
By the same token if $P$ maximizes $U$ subject to a marginal constraint $P|_{\xi}=\nu$, then $Y$ should be additively separable. More on this later.
@@
Schrödinger Potentials
The problem faced by the DM can be separated into two problems: i) maximizing the objective eq.˜1 subject to a further marginal constraint $P|_{\xi}=\nu$ and ii) solving for the optimal marginal $\nu$. We observe here that the former is an instance of the entropic optimal transport problem. The constrained optimal marginal is called the Schrödinger bridge from $\nu$ to $\mu$.
In lemma˜8, we established that if $P$ maximizes $U$ subject to feasibility and Bayes plausibility, then $P|_{\xi}\sim\phi$ and $dP(\xi)/d\phi(\xi)$ is upper bounded and lower bounded away from zero. So in this section, we assume $P|_{\xi}$ is restricted to a marginal $\nu$ for which $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from zero.
Define ${\cal Q}(\nu)=\{P\in\Delta(X\times\Theta):P|_{\xi}=\nu,P|_{\theta}=\mu\}$. We let
$V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$
denote the value of imposing the constraint $P|_{\xi}=\nu$.
@@
$-V(\nu)=\inf_{P\in{\cal Q}(\nu)}\int c(\xi,\theta)\>dP-D_{KL}(P\|\nu\otimes\mu)$ (4)
where $c(\xi,\theta)=-{u(\xi,\theta)}+\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)\in L^{\infty}(\nu\otimes\mu)$.
This is called the Entropic Optimal Transport (EOT) problem. [^3] For simplicity, we may assume $c(\xi,\theta)$ is totally bounded, since the worse case scenario choice of $\nu$ can be set to $\phi$, in which case c=uc=-u, which we assumed to be totally bounded. The following theorems concern properties of EOT and can be found in [6], up to notational changes.
@@
If $P$ solves eq.˜4, then there exist functions $a_{\nu}\in L^{1}(\nu),-b_{\nu}\in L^{1}(\mu)$ such that
@@
These are called the Schrödinger potentials. Conversely, if there exists potentials such that eq.˜5 holds, then $P$ must be optimal. Moreover, potentials are unique up to translation. [^4] That is, for any two pairs of potentials $(a_{\nu},b_{\nu})$ and $(a_{\nu}^{\prime},b_{\nu}^{\prime})$, we have $a_{\nu}-a^{\prime}_{\nu}=b_{\nu}^{\prime}-b_{\nu}$.
@@
If $P$ solves eq.˜4 then $Y(\xi,\theta;P)=a_{\nu}(\xi)+b_{\nu}(\theta)$ and so $U(P)=E_{\nu}[a_{\nu}(\xi)]+E_{\mu}[b_{\nu}(\theta)]$
@@
$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}}\>d\nu(\xi)}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{b_{\nu}(\theta)}}\>d\mu(\theta)}\>d\nu(\xi)$
@@
Note that because $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from 0, by Hölder’s equality with coefficients $p=1$ and $q=\infty$, we can interchange $a\in L^{1}(\nu)$ with $a\in L^{1}(\phi)$.
Corollary 11.1.
Consider a path $\nu_{\varepsilon}$ such that at $\varepsilon=0$, $\varepsilon\mapsto\frac{d\nu_{\varepsilon}(\xi)}{d\phi(\xi)}$ is continuous for $\phi$-a.e. $\xi$, and $d\nu_{\varepsilon}/d\phi$ is uniformly bounded away from 0 and $\infty$ on $\varepsilon\in(-r,r)$. Then $\varepsilon\mapsto V(\nu_{\varepsilon})$ is continuous at 0.
In the discrete case, the Schrödinger potentials can be viewed as the Lagrange multipliers for the marginal constraints. In other words, $a_{\nu}(x)$ can be viewed as the rate of increase in $V$ when adding an infinitessimal mass of $x$ to $\nu$. This is because
$V(\nu)=\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
so in the discrete case, applying the Envelope theorem and normalizing $\int a_{\nu}\>d\nu=0$ gives
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}V\bigg{(}(1-\varepsilon)\nu+\varepsilon\delta_{x}\bigg{)}=a_{\nu}(x)$
The continuous case is more involved and discussion is deferred to the appendix.
@@
If one writes out the objective – a function of the joint $P$ – swaps the log with the first integral, one obtains a function of the marginal $P|_{\xi}$. When $\alpha\in(0,1)$ this function is strictly convex, and, by Jensen’s Inequality, upper bounds $V$. We show that their maxima coincide, meaning that solving for the optimal marginal is a relatively simple convex optimization problem.
Define $f(\nu)=E_{\theta\sim\mu}\left[\log\left(E_{\xi\sim\phi}\left[e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}\right]\right)\right]\equiv E_{\theta\sim\mu}[\log(Z(\theta;\nu)]$
@@
If $P$ maximizes $U$ then $f(P|_{\xi})=U(P)$.
@@
The maximum of $f$ and $U$ coincide. That is, if $\nu$ maximizes $f$, then there exists a $P$ with $P|_{\xi}=\nu$ which maximizes $U$.

### v2

@@ Introduction
Example: state-action with costlessly observed $\theta$. Again, consider a state-action microfoundation in which the DM knows their state $\theta$ a priori. The DM has an action set $A=\{1,...,J\}$ consisting of $J$ products with $J$ characteristics. The DM knows what the characteristics are, but does not know how they are permuted among the $J$ actions. Let $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$ denote these characteristics. Denote $\xi$ with the characteristic of the product he eventually chooses. Given $\theta$, the mutual information $I_{P}(j,\boldsymbol{\xi})$ between action $j\in A$ and $\boldsymbol{\xi}$ is
$\sum_{j\in A}D_{KL}(P(\boldsymbol{\xi}|j,\theta)\|P(\boldsymbol{\xi}))\cdot P(j)=D_{KL}(P(\boldsymbol{\xi}|j=1,\theta)\|P(\boldsymbol{\xi}))=D_{KL}(P(\xi_{1}|j=1,\theta)\|P(\xi_{1}))$
@@
which is equal to $D_{KL}(P(\xi|\theta)\|\phi(\xi))$. The third equality follows from the observation that choosing $j=1$ provides only information about what $\xi_{1}$ and is not (meaning it provides no information about $(\xi_{2},...,\xi_{J})$ beyond the fact that they are not $\xi_{1}$). Thus, this corresponds to the state-characteristic model with $\alpha=1$ and $\phi(\xi)=1/J$.
@@
Connection to Maxwell-Boltzmann statistics. When $\alpha=1$, the law of iterated expectations yields
@@
$\frac{dP(\xi,\theta)}{d(P_{\xi}\otimes\mu)(\xi,\theta)}=\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}<\infty$ $\frac{dP(\xi)}{d\phi(\xi)}<\infty$ $\frac{dP(\theta)}{d\mu(\theta)}=1$ (4)
@@
If $P_{\xi}\ll\phi$ and $P_{\theta}=\mu$, then $P(\cdot|\theta)\ll\phi$ $\mu$-a.s. and $P\ll\phi\otimes\mu$.
@@
Let $P_{\xi}\ll\phi$ and $P_{\theta}=\mu$. There exist valid densities such that eq.˜4 is satisfied.
@@
Because results have been established for $\alpha\in\{0,1\}$, we focus on the case where $\alpha\in(0,1)$. We normalize $\lambda$ to $1$, since maximizing $U(P)$ is the same as maximizing $U(P)/\lambda$. To obtain results for $\lambda\neq 1$, one can simply take all results herein in which $u(\xi,\theta)$ and associated additive terms appear and divide them by $\lambda$.
For clarity of exposition, it helps to write out $U(P)$ as a single integral
$U(P)=\int Y(\xi,\theta;P)\>dP$
@@
If $P$ solves the objective eq.˜3 then
@@
If $P$ solves the objective eq.˜3 then in the discrete case,
@@
The optimal marginal $P_{\xi}$ is unique, and therefore the optimal CCP is unique.
@@
If $P$ maximizes $U$, then $P_{\xi}\ll\phi\ll P_{\xi}$.
These follow from corollaries˜7.1, 8 and 16.1. In the discrete case, these can be derived by taking Lagrangians.
@@
$\frac{P(\xi|\theta)}{P^{\prime}(\xi|\theta)}=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}}{\phi^{\prime}(\xi)^{\alpha}P^{\prime}(\xi)^{1-\alpha}}\frac{Z^{\prime}(\theta;P^{\prime}_{X})}{Z(\theta;P_{\xi})}$
@@
First-Step Orthogonality and the Gibbs Property
We begin our mathematical results sections by highlighting a few necessary conditions for optimality. First-step orthogonality is a property of moments of the form $\int f_{P}(\omega)\>dP(\omega)$, where the integrand $f$ depends on the probability with respect to which we are integrating. Loosely speaking, a moment is said to be first-step orthogonal [^3] This term is borrowed from the double machine learning literature. perturbation in $P$ locally affects the moment only through the re-weighting $dP$. It constitutes a sufficient condition for the Gibbs property and additive separability. It is a relatively unique property of entropy-based costs which essentially comes from the fact that $\frac{d}{dx}\log x=1/x$.
@@
By the same token if $P$ maximizes $U$ subject to a marginal constraint $P_{\xi}=\nu$, then $Y$ should be additively separable. More on this later.
@@
Lemma˜8 is a crucial lemma because it guarantees that the optimal marginal $P_{\xi}$ lies on the interior, so it is equally crucial to transparently pinpoint its origin: effectively, it arises from an “Inada condition” on the functional form of the cost of information gain from $P_{\xi}$ to $\phi$. To highlight this, consider the discrete case, where the information cost $\kappa_{\alpha}$ can be written as
$\kappa_{\alpha}$ $=\alpha D_{KL}(P_{\xi}\|\phi)+I_{P}(\xi,\theta)$
$=\left[\alpha\sum_{\xi\in X}\left(\frac{P(\xi)}{\phi(\xi)}\right)\log\left(\frac{P(\xi)}{\phi(\xi)}\right)\ \phi(\xi)\right]+I_{P}(\xi,\theta)$
$\equiv\left[\alpha\sum_{\xi\in X}f(\xi)\log(f(\xi))\ \phi(\xi)\right]+I_{P}(\xi,\theta)$
where $f=P_{\xi}/\phi$ is the likelihood ratio/density. For $\alpha>0$ and $\xi$ such that $\phi(\xi)>0$, as $f(\xi)\to 0$, $\frac{\partial\kappa}{df(\xi)}\to$-$\infty$; that is, the marginal cost of reducing $f(\xi)\equiv P(\xi)/\phi(\xi)$ increases to infinity.
In words, the result that the DM never completely zeroes out the possibility of choosing a lemon ($P(\text{lemon})$) when such lemons exist ($\phi(\text{lemon})>0$) in the market comes from the assumption that as the chance of choosing a lemon goes to 0, it becomes arbitrarily costly to further reduce those chances. This is mathematically embodied in this model by the Inada property of $x\log x$.
Rational Inattention as Entropic Optimal Transport
The problem faced by the DM can be separated into two problems: i) maximizing the objective eq.˜3 subject to an additional marginal constraint $P_{\xi}=\nu$ and ii) solving for the optimal marginal $\nu$. We observe here that the former is an instance of the entropic optimal transport problem. The constrained optimal marginal is called the Schrödinger bridge from $\nu$ to $\mu$.
Define $\Pi(\nu,\mu)=\{P\in\Delta(X\times\Theta):P_{\xi}=\nu,P_{\theta}=\mu\}$ to be the set of all joint distributions with marginal restrictions $\nu$, $\mu$. Then, we define the value of imposing the second constraint $\nu$ to be
$V(\nu)=\sup_{P\in\Pi(\nu,\mu)}U(P)$
denote the value of imposing the constraint $P_{\xi}=\nu$.
@@
$-V(\nu)=\inf_{P\in\Pi(\nu,\mu)}\int\underbrace{-u(\xi,\theta)+\alpha D_{KL}(\nu\|\phi)}_{c(\xi,\theta)}\>dP+D_{KL}(P\|\nu\otimes\mu)$ (6)
where $c$ is uniformly bounded from below by assumption. This is the Entropic Optimal Transport (EOT) problem. The following theorems apply the key properties of EOT to the state-characteristic model and can be found in [nutz2021introduction], up to notational changes.
Theorem 9.
For each $\nu$, there exists a unique Schrödinger bridge $P\in\Pi(\nu,\mu)$ that attains the infimum in eq.˜6.
@@
Let $P$ be the Schrödinger bridge that solves eq.˜6. There exist functions $a_{\nu}\in L^{1}(\nu),-b_{\nu}\in L^{1}(\mu)$ such that
@@
These are called the Schrödinger potentials. Conversely, if there exist potentials such that eq.˜7 holds, then $P$ must be optimal. Moreover, potentials are unique up to translation. [^4] That is, for any two pairs of potentials $(a_{\nu},b_{\nu})$ and $(a_{\nu}^{\prime},b_{\nu}^{\prime})$, we have $a_{\nu}-a^{\prime}_{\nu}=b_{\nu}^{\prime}-b_{\nu}$.
The statement above says that $a_{\nu}$ is in $L^{1}(\nu)$; but because $\nu$ is endogenous, we would much rather have a statement that says $a_{\nu}\in L^{1}(\phi)$. We established in lemma˜8 that if $\alpha>0$, then $\frac{d\nu}{d\phi}$ is bounded away from 0 and $\infty$. Thus, a stronger statement holds:
@@
Suppose $\frac{d\nu}{d\phi}$ is bounded away from 0 and $\infty$. Then, $f\in L^{1}(\nu)\Leftrightarrow f\in L^{1}(\phi)$. $P$ solves eq.˜6 if and only if there exist functions $a_{\nu}\in L^{1}(\phi),-b_{\nu}\in L^{1}(\mu)$ satisfying eq.˜7.
Proof.
Let $f\in L^{1}(\nu)$. Then
$\bigg\|\frac{d\nu}{d\phi}\bigg\|_{\infty}\int|f|\ d\nu\geq\int|f|\ \frac{d\nu}{d\phi}\ d\phi=\int|f|\ d\phi$
Conversely, if $f\in L^{1}(\phi)$ then
$\bigg\|\frac{d\phi}{d\nu}\bigg\|_{\infty}\int|f|\ d\phi\geq\int|f|\ \frac{d\phi}{d\nu}\ d\nu=\int|f|\ d\nu$
A trivial consequence is additive separability of the integrand when $P$ is optimal.
Corollary 10.2.
If $P$ solves eq.˜6 then $Y(\xi,\theta;P)=a_{\nu}(\xi)+b_{\nu}(\theta)$ and so $U(P)=E_{\nu}[a_{\nu}(\xi)]+E_{\mu}[b_{\nu}(\theta)]$
Two results from EOT theory prove useful for studying this system. The first are the Schrödinger Equations which characterise the relationship between the Schrödinger potentials. The second is the duality formula for EOT.
@@
$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi^{\prime},\theta)}}{e^{a_{\nu}(\xi^{\prime})}\left(\frac{d\nu(\xi^{\prime})}{d\phi(\xi^{\prime})}\right)^{\alpha}}\>d\nu(\xi)}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta^{\prime})}}{e^{b_{\nu}(\theta^{\prime})}}\>d\mu(\theta^{\prime})}\>d\nu(\xi)$
If $P$ is optimal and $\nu=P_{\xi}$ is the marginal, then the potential $a_{\nu}(\xi)$ should be constant, since the marginal for $X$ is a choice variable. After normalizing $a_{\nu}=0$, it should be obvious that the potential $b_{\nu}(\theta)$ equals the log of the partition function:
$b_{\nu}(\theta)=\log(Z(\theta;P))$
@@
If one writes out the objective – a function of the joint $P$ – swaps the log with the first integral, one obtains a function of the marginal. When $\alpha\in(0,1)$ this function is strictly concave, and, by Jensen’s Inequality, upper bounds $V$. We show that their maxima coincide, meaning that solving for the optimal marginal is a relatively simple optimization problem.
Define
$f(\nu)=E_{\theta\sim\mu}\left[\log\left(E_{\xi\sim\phi}\left[e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}\right]\right)\right]\equiv E_{\theta\sim\mu}[\log(Z(\theta;\nu))]$
and, as before,
$V(\nu)\equiv\sup_{P\in\Pi(\nu,\mu)}U(P)=\sup_{P\in\Pi(\nu,\mu)}E_{(\xi,\theta)\sim P}\left[\log\left(e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\left(\frac{dP(\xi,\theta)}{d(\nu\otimes\mu)(\xi,\theta)}\right)^{-1}\right)\right]$
The primary claim of this section is that
$\sup_{\nu\ll\phi}f(\nu)=\sup_{\nu\ll\phi}V(\nu)$
The secondary claim is that $f$ is strictly concave. These two claims imply that it suffices to maximize the strictly concave function $f(\nu)$.
Theorem˜13 is an application of Jensen’s inequality, establishing that $f\geq V$. Theorem˜14 establishes that $f$ strictly concave and hence is uniquely maximized. It follows that to show $V$ is uniquely maximized, it suffices to show that the maximum of $f$ and $V$ coincide, as shown in theorem˜16. The immediate corollary˜16.1 is that $V$ is uniquely maximized, and therefore $U$ is uniquely maximized.
@@
If $P$ maximizes $U$ subject to Bayes plausibility, then $f(P_{\xi})=U(P)$. Hence, $f(P_{\xi})=\sup_{\nu\ll\phi}V(\nu)$
@@
The maximum of $f$ and $V$ coincide. That is, if $\nu$ maximizes $f$, then there exists a $P$ with $P_{\xi}=\nu$ which maximizes $U$.
Because each marginal $\nu$ admits a unique Schrödinger bridge $P\in\Pi(\nu,\mu)$, we arrive at our main result:
@@
Discussion
Schrödinger potentials can be interpreted as the infinite-dimensional analog to the Lagrange multiplier. The Schrödinger potential $a_{\nu}(\xi)$ represents a shadow “probability price” on $\nu(\xi)$; it follows that if $\nu$ is optimal, then $a_{\nu}(\xi)$ should be constant a.e. Plugging this fact into the Schrödinger equations, it follows that the optimal marginal $\nu$ satisfies
$\int\frac{e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}}{\int e^{u(\xi^{\prime},\theta)}\left(\frac{d\nu(\xi^{\prime})}{d\phi(\xi^{\prime})}\right)^{1-\alpha}\>d\phi(\xi^{\prime})}\>-1\ d\mu(\theta)=0$ (8)
This is a state-characteristic analog of the necessary and sufficient condition derived by [caplin2019rational]. [^5] Recall that theirs involved an inequality because when $\alpha=0$, there can be “corner solutions.” From staring at the definition of $f$, one can glean that eq.˜8 is a first-order condition for $f$ to be maximized on the interior. [^6] If one needs any convincing, set $X$ to be discrete, and write out the first-order condition corresponding to a Gateaux derivative perturbation in the direction of a Dirac probability mass $\delta_{x}$. As theorem˜14 showed, $f$ is strictly concave on a convex set, implying a unique point satisfying this first-order condition, if it exists. It is then left to show that the maximum of $f$ likewise does not lie on the boundary.
Computationally, eq.˜8 implies that finding the optimal $\nu$ amounts to finding the fixed-point of
$Tg(\xi)=\int\frac{e^{u(\xi,\theta)}g(\xi)^{1-\alpha}}{\int e^{u(\xi^{\prime},\theta)}g(\xi^{\prime})^{1-\alpha}d\phi(\xi^{\prime})}\>d\mu(\theta)$
In the model, our goal was to maximize $U$ subject to one marginal constraint $P_{\theta}=\mu$. But since we have taken this unusual approach of first maximizing subject to two constraints $P\in\Pi(\nu,\mu)$ and then optimizing the first constraint $\nu$, it makes ample sense to consider the marginal benefit of relaxing our constraints.
In the discrete case, the Schrödinger potentials can be directly interpreted as the Lagrange multipliers for the marginal constraints. In other words, $a_{\nu}(x)$ can be viewed as the rate of increase in $V$ when transfering an infinitessimal mass away from all other elements of $X$ to $\nu(x)$. This is because
$V(\nu)=\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
via duality, so in the discrete case, applying the Envelope theorem and normalizing $\int a_{\nu}\>d\nu=0$ gives
$\frac{d}{d\varepsilon}\bigg|_{\varepsilon=0}V\bigg((1-\varepsilon)\nu+\varepsilon\delta_{x}\bigg)=a_{\nu}(x)-\displaystyle\int a_{\nu}\>d\nu=a_{\nu}(x)$
where $\delta_{x}$ is the Dirac probability at $x$. The continuous case is more involved and discussion is deferred to the appendix. For now, let us assume the premise is true.
To consider the marginal benefit of relaxing a constraint, consider a concrete example of a DM choosing an insurance policy. Utility depends on consumption, which at time 0 is given as $c_{0}$ and at time 1 is given as $c_{1}(\xi,\theta)$ – depending on insurance contract $\xi$ and state $\theta$. Suppose we restricted the DM to choosing information policies $P\in\Pi(\nu,\theta)$, where $\nu$ is sub-optimal. Then the marginal benefit in utils of shifting $\nu$ infinitessimally towards $\delta_{x}$ is $a_{\nu}(x)$. Using logarithmic utility (which makes sense since we assume entropic costs) for time 0 consumption, i.e.,
$V_{0}(\nu)=\log c_{0}-\alpha D_{KL}(\nu\|\mu)+\sup_{P\in\Pi(\nu,\mu)}\iint u(c_{1}(\xi,\theta))\>dP(\xi,\theta)-I_{P}(\xi,\theta)$
we could then use marginal utility of consumption $\partial V_{0}/\partial c_{0}=1/c_{0}$ as a welfare numeraire. $a_{\nu}(x)\cdot c_{0}$ can be interpreted as the “probability price” in terms of time 0 consumption – the willingness to pay – to shift $\nu$ marginally towards $\delta_{x}$.
Likewise, if we instead normalized $b_{\nu}$, then $b_{\nu}(\theta)$ can be viewed as the “probability price” of changing the underlying probability distribution over states of the world to add additional probability to state $\theta$. Because $\mu$ is generally not optimal – the DM takes it as given – this is more of a hypothetical. However, it could be useful for considering the welfare effect of actual shifts in probability – for example, the welfare change from decreasing the probability of a certain risk, given that consumers are rationally inattentively insured.
Computation
We turn to the computational problem of finding the optimal joint distribution $P$ which solves eq.˜3 when $X$ and $\Theta$ are finite sets. By eq.˜5, it suffices to solve for the optimal marginal $P_{\xi}$ to obtain the optimal conditional choice probability. We suggest three algorithms for doing so.
Fixed-Point Iteration. [caplin2019rational] suggest the Blahut-Arimoto (BA) algorithm as a method for computing the optimal value. The direct analogue of BA is
$g_{t+1}(\xi)=[Tg_{t}](\xi)=\int\frac{e^{u(\xi,\theta)}g_{t}(\xi)^{1-\alpha}}{\int e^{u(\xi^{\prime},\theta)}g_{t}(\xi^{\prime})^{1-\alpha}d\phi(\xi^{\prime})}\>d\mu(\theta)$ $g_{t}(\xi)=\frac{d\nu_{t}(\xi)}{d\phi(\xi)}$
and takes advantage of the fact that $g$ corresponds to the optimal $\nu$ if and only if $Tg=g$. Fixed-point iteration converges nicely for the majority of reasonably conceivable problems.
Convex optimization. An alternative to fixed-point iteration is convex optimization of f-f. The results establish f-f is a strictly convex function over the convex set $\Delta X$ which is known to be minimized on $\text{int}(\Delta X)$, and that the $\nu$ which minimizes f-f is the $\nu$ which maximizes $V$.
Numerical convex optimization methods like mirror ascent work reasonably well but can scale poorly. The main pitfall of this is that if the optimal $\nu$ lies near the boundary, the slope between $\nu$ and the boundary is extremely steep.
Sinkhorn’s Algorithm. For discrete $\xi,\theta$, the problem of finding Schrödinger potentials such that eq.˜7 holds is equivalent to the matrix scaling problem. Formally, we can re-write eq.˜7 in the discrete case as
$P(\xi,\theta)=e^{-a_{\nu}(\xi)}\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]e^{-b_{\nu}(\theta)}$ (9)
If we think of the middle term as a $|X|\times|\Theta|$ matrix $A$ and $P$ as a $|X|\times|\Theta|$ matrix such that the marginals are $\nu$ and $\mu$ respectively, we can re-formulate the problem as follows: given a matrix A and a pair of probability vectors $\boldsymbol{\nu}$, $\boldsymbol{\mu}$, find diagonal matrices $\textbf{D}_{1},\textbf{D}_{2}$ such that for $\textbf{P}:=\textbf{D}_{1}\textbf{A}\textbf{D}_{2}$ is a probability matrix whose columns sum to $\sum_{\xi}P_{\xi\theta}=\mu_{\theta}$ and rows sum to $\sum_{\theta}P_{\xi\theta}=\nu_{\xi}$. For our purposes, A is the matrix whose $(\xi,\theta)$-entry is $\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]$.
The matrix scaling problem has a known solution: Sinkhorn’s algorithm. For a given A, Sinkhorn’s algorithm solves for the optimal potentials $a_{\nu}$, $b_{\nu}$, which allow us to back out the optimal $P\in\Pi(\nu,\mu)$ via eq.˜9.
Theorem 17 (Sinkhorn’s algorithm).
[nutz2021introduction, Theorem 6.15] Consider the EOT problem. Set $a_{t}=0$. For $t\geq 0$, set
$a_{t+1}(\xi)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)-\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)$
Then:
  1. 1.
  2. 2.
    $D_{KL}(P\|P_{t})\to 0$
  3. 3.
    $D_{KL}(P_{t}\|R)\to D_{KL}(P\|R)$
  4. 4.
    $P_{t}\to P$ in total variation
Although Sinkhorn’s algorithm allows us to find the Schrödinger potentials for a given $\nu$, it does not tell us what the optimal $\nu$ should be.
However, observe that the Schrödinger potential $a_{t}(x)$, after normalization, tells us how much $V$ improves when we update $\nu$ to add mass to $x$. It can be thought of as the “direction of steepest ascent,” and so it makes sense to integrate a “gradient ascent” step into the algorithm which does not even require computing a gradient (since we already have $a_{t}$ from Sinkhorn iteration). One such procedure is given by
$b_{t}(\theta)=\log\left(\int e^{u(\xi,\theta)-a_{t}(\xi)}\>d\phi^{\alpha}\>d\nu_{t}^{1-\alpha}\right)$
$a_{t+1}(\xi)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)-\alpha\log\left(\frac{d\nu_{t}(\xi)}{d\phi(\xi)}\right)$
$d\nu_{t+1}(\xi)=\frac{d\nu_{t}(\xi)\exp\left(\eta_{t+1}a_{t+1}(\xi)\right)}{\int\exp\left(\eta_{t+1}a_{t+1}(\xi^{\prime})\right)d\nu_{t}(\xi^{\prime})}$
where $\eta_{t}$ is tunable. The update on $\nu_{t+1}$ is equivalent to
$\log\left(\frac{d\nu_{t+1}(\xi)}{d\phi(\xi)}\right)\propto\log\left(\frac{d\nu_{t}(\xi)}{d\phi(\xi)}\right)+\eta_{t+1}\left(a_{t+1}(\xi)-E_{\phi}[a_{t+1}]\right)$
The above-specified algorithm updates $\nu$ every time $a$ and $b$ are updated, but this can also be tuned (i.e. to allow for several updates of $a$ and $b$ for every update of $\nu$).
This procedure can be faster than the other methods, especially for large $|X|\times|\Theta|$. Generally, it works better when the optimal $\nu$ lies farther from the boundary, which is typically the case when the cost of acquiring information about $\xi$ – i.e. $\alpha\lambda$ – is not too low.